An aerobatic aircraft is performing circular loops at the speed of sound, covering spectacularly 340 meters every single second in a flight exhibition. All participants have been warned that they keep their acceleration below 8 times that of the gravitational acceleration to avoid getting dizzy. What is the diameter of the smallest loop that these participants can safely do? Here you may use 10 m/s/s as the gravitational acceleration.


1
Expert's answer
2021-05-27T17:47:57-0400

Solution.

"v=340m\/s;"

"a=\dfrac{g}{8}m/s^2;"

"g=10m\/s^2;"

"a=\dfrac{v^2}{r}\implies r=\dfrac{v^2}{a};"

"r=\dfrac{(340m/s)^2}{\dfrac{10}{8}m/s^2}=92480m;"

"d=2r=2\sdot 92480m=184960m;"

Answer: "d=184960m."